A group of 7 adults and 10 children went swimming. The entrance fee of an adult is twice the amount for the children. The group paid a total of Php. 1,440. How much was the entrance fee of an adult?
Answer by ikleyn(52788)

.
7x + 10*(0.5x) = 1440


7x + 5x = 1440


12x = 1440


x = 1440/12 = 120.

Answer.   Php 120.